# Can repeating and non-repeating FRBs be drawn from the same population?

### v1

@@ Abstract
Do all Fast Radio Burst (FRB) sources repeat? We present evidence that FRB sources follow a Zipf-like distribution, in which the number density of sources is approximately inversely proportional to their burst rate above a fixed energy threshold—even though both the burst rate and number density span many orders of magnitude individually. We introduce a model-independent framework that predicts the distribution of observed fluences and distances, and repetition rates of an FRB population based on an assumed burst rate distribution per source. Using parameters derived directly from observations, this framework simultaneously explains several key features of the FRB population: (i) The observed ratio of repeaters to apparent non-repeaters; (ii) The much lower ratio of apparent non-repeaters to the total number of Soft Gamma Repeater (SGR) sources within the observable Universe; And (iii) the slightly smaller average distances of known repeaters compared to non-repeaters. We further explore how survey parameters, such as radio sensitivity and observation time, influence these statistics. Notably, we find that the fraction of repeaters rises only mildly with improved sensitivity or longer exposure. This weak dependence could be misinterpreted as evidence that not all FRBs repeat. Overall, our results support the idea that a single population—likely magnetars—can account for the full observed diversity of FRB activity, from very inactive FRB sources like SGR 1935+2154 to the most active repeaters.
@@ 1 Introduction
Fast radio bursts (FRBs) exhibit a striking diversity in their observable properties. The isotropic-equivalent energies of extragalactic FRBs with known redshifts span over nine orders of magnitude - from as low as $\sim 6\times 10^{23}\mbox{erg Hz}^{-1}$ in FRB 20200120E (Nimmo et al., 2023) to as high as $\gtrsim 6\times 10^{32}\mbox{erg Hz}^{-1}$ in FRB 20220610A (Ryder et al., 2023). Even individual repeaters can show large variability: for example, bursts from FRB 20201124A cover about five orders of magnitude in energy (Zhang et al., 2022; Kumar et al., 2022). FRB sources also differ greatly in their apparent activity rate (Margalit et al., 2020). Some, like FRBs 20121102A and 20201124A, have been observed to emit dozens of bursts in a single hour (Li et al., 2021; Zhang et al., 2022) while others - such as the Galactic magnetar SGR 1935+2154 - have only produced a single energetic burst with in an observed period of over several years. As discussed in §3, this contrast becomes even more pronounced when comparing sources at similar energies, due to the steep energy dependence of burst rates. A third axis of variation is the inferred number density of sources. For instance, Lu et al. (2022) showed that the volumetric density of highly active extragalactic FRBs like 20180916B and 20121102A is orders of magnitude lower than that of low-activity sources such as SGR 1935+2154 or FRB 20200120E in the nearby M81 galaxy. Given these dramatic differences - in energy output, repetition rate, and source density - it remains an open question whether all FRBs can be unified under a single physical framework, and if so, what underlying parameters account for their apparent diversity.
@@
Perhaps the most striking of the apparent dichotomies, is the observational separation between sources seen only once thus far (“non-repeaters") and those seen multiple times (“repeaters"). The most reliable way to estimate the intrinsic fraction of repeating sources is to consider a blind survey mission such as CHIME. In the first 20 months of its operation, CHIME detected 46 repeating sources (and 14 more candidates), consisting a fraction of $2.6_{-2.6}^{+2.9}\%$ of sources in their sample (Chime/Frb Collaboration et al., 2023). The lack of bimodality in burst rates per source between the population of repeaters and (as of yet) non-repeaters detected by CHIME (with observed rates varying continuously from a few per hour to $\lesssim 10^{-3}\mbox{ hr}^{-1}$), suggests all FRB sources might repeat, and that the low observed fraction is just a reflection of the current sensitivity. The finding that repeaters in this sample have, on average, slightly lower extragalactic DMs than non-repeaters ($\sim 4300\mbox{ pc cm}^{-3}$ as compared with $\sim 600\mbox{ pc cm}^{-3}$), could simply be a selection effect - repeaters are by the nature of their identified repetition, likely to be slightly closer than non-repeaters. This does not require intrinsically distinct populations. Similarly, while there are some small differences between repeater and non-repeater host galaxies (the former being slightly less massive, and having lower metalicity, Sharma et al. 2024), these differences are not statistically meaningful and might in addition be subject to further selection biases.
@@
A different way to asses the possibility of a shared origin of repeaters and non-repeaters, is to monitor over a long time known, highly active, repeater sources. The energy distribution per source is found (at least by some studies) to be consistent with the energy distribution of non-repeaters (Wang & Yu, 2017; Zhang et al., 2021). Kirsten et al. (2024) have observed FRB 20201124A for over 2000 hours, using 25-32 m class telescopes. They have found that the slope of the burst energy distribution varies both as a function of time (sporadically) and energy (becoming shallower at higher energies). The higher energy slope is consistent with that of non-repeating sources and the time variation helps explain the apparent large range of repeating rates between sources. Similar results (utilizing 1500 hours of observations) were found for another repeater, FRB 20220912A (Ould-Boukattine et al., 2024).
In this paper, we set out to explore whether the activity rates, source densities, and energies of FRB repeaters / non-repeaters can originate from the same underlying population. Our approach is primarily analytic (backed by numerical calculations) and is independent of the nature of FRB sources. We begin, in §2 by exploring the simplest hypothesis supporting a unified FRB origin: that there is a single source type powering all observed FRBs. Then, in §3 (and §B) we expand the model to allow for variable intrinsic rates and typical burst energies. The results are then compared with observations in §4 and implications for magnetar sources are discussed in §5. We conclude in §6.
@@ 2 Single source rate
for $E_{\rm min}<E_{\nu}<E_{\rm c}$ and $\gamma>1$, where $E_{\nu}$ is the energy per unit frequency. $E_{*}$ is the energy used for the normalization of the rate per source. Throughout the majority of this work we choose $E_{\rm*}=E_{\rm c}$ (an exception to this is discussed in §B).
@@
At the simplest level, a large sky FRB survey can be described by its limiting (spectral) fluence $\Phi_{\rm lim}$ and by the time, $T$, it spends observing each point on the sky  [^3] We first consider a uniform value of $T$ across the sky; we provide a generalized treatment in §A. During the data collection period for the 1st CHIME catalog, the median exposure time was $\sim\!20$ hrs (CHIME/FRB Collaboration et al., 2021), which we adopt as a reference value when using a fixed $T$.. While lower energy bursts are more common per source (Eq. 1), such bursts cannot be detected to large distances by a fluence-limited survey. At the same time the number of sources increases rapidly with distance. The result is a competition between the detectability of low energy and close-by sources and high energy far-away sources. As detailed below (see also Lu & Piro 2019), for $\gamma<2.5$ the competition is won by the far-away sources. That is, most detected bursts are from large cosmological distances. However, as we describe next, this conclusion does not necessarily hold for active repeaters. This is because, if a single burst energy / distance dominates the entire sample, then this corresponds to a single mean number of repetitions per source during the observed period, $\lambda_{E_{\nu}}=\mathcal{R}(>E_{\nu})T$444a verification of our results using all relevant cosmological corrections is presented in §A.. Assuming Poisson statistics, the probability that a source will repeat $k$ times during $T$ is $P(k)=\lambda^{k}e^{-\lambda}/k!$  [^5] Several studies reported non-Poissonian burst arrival time statistics from prolific repeaters, with bursts tending to cluster in time (Oppermann et al., 2018; Wang et al., 2018; Oostrum et al., 2020). However, Cruces et al. (2021); Jahns et al. (2023) showed that once short separations (on the order of tens of ms, likely corresponding to non-independent sub-bursts) are excluded and periodic activity is accounted for, the remaining distribution is consistent with a Poisson process. Since deviations from Poissonian behavior would require additional model parameters, we adopt a Poissonian model here to avoid overfitting. This assumption can be revisited if future observations provide strong evidence for non-Poissonian statistics.. Therefore, $\lambda/2$ is approximately the ratio between repeaters (sources observed at least twice) and non-repeaters and calibration with the observed ratio leads to $\lambda\approx 0.05$. $P(k)$ becomes vanishingly small for large $k$. For example $P(k\geq 10)/P(k\geq 1)\approx\lambda^{10}/10!\sim 4\times 10^{-20}$. The fact that CHIME has detected multiple sources with more than 10 repetitions clearly means that such a picture is too simplistic and that the distance and / or intrinsic rate per source must be varying for sources detected multiple times. The first point is addressed below and the second in §3.
We define $r_{z}$ as the proper distance up to $z\approx 2$. Beyond $r_{z}$ the number of sources increases much more slowly, due to a significant departure from Euclidean geometry (as well as a likely intrinsic reduction in the source density, at eras before the peak of star formation history). The combination of $r_{z}$ and $\Phi_{\rm lim}$ yields the typical energy of detectable non-repeater bursts $E_{\rm lim}\equiv 4\pi r_{z}^{2}\Phi_{\rm lim}$ (those that are close to the threshold fluence and arise from distant sources), which corresponds to $\lambda_{\rm lim}\equiv\lambda_{\rm E_{\rm lim}}$. The discussion above suggests that $\lambda_{\rm lim}\approx 0.05$ and
@@
$\mathcal{R}(>E_{\rm lim})\!\approx\!3\times 10^{-3}\left(\frac{20\mbox{ hr}}{T}\right)\left(\frac{\lambda_{\rm lim}}{0.05}\right)\mbox{ hr}^{-1}$
@@
As $\lambda_{E_{\nu}}$ increases with decreasing $E_{\nu}$, one can define a critical energy, $E_{\rm all,k}$ such that $P(>E_{\rm all,k})\approx 1$ (this energy approximately satisfies $\lambda(>E_{\rm all,k})=k$). This leads to
@@
where $\lambda_{\rm c}\!\equiv\!\lambda_{E_{\rm c}}$ and the last transition necessarily holds for $\lambda_{\rm lim}\!<\!1$. One can equivalently define critical fluences $\Phi_{\rm c}\!=\!E_{\rm c}/(4\pi r_{z}^{2})$ and $\Phi_{\rm all,k}\!=\!E_{\rm all,k}/(4\pi r_{z}^{2})\!\approx\!(\lambda_{\rm c}/k)^{1\over\gamma-1}\Phi_{\rm c}$. Below $\Phi_{\rm all,k}$, roughly all sources in the Universe are observed by the survey $k$ times. Taking $E_{\rm c}\!\approx\!3\!\times\!10^{33}\mbox{erg Hz}^{-1},\gamma\!\approx\!1.7$ as inferred from FRB energy distribution studies calibrated by observations (Lu & Piro, 2019; James et al., 2022; Hashimoto et al., 2022; Shin et al., 2023; Lin & Zou, 2024; Gupta et al., 2025), we find $\lambda_{\rm c}\!\approx\!7\!\times\!10^{-3},\Phi_{\rm c}\!\approx\!100\mbox{Jy ms},\Phi_{\rm all,1}\!\approx\!0.07\mbox{Jy ms}$, so we expect the ordering $\Phi_{\rm all,1}\!\ll\!\Phi_{\rm lim}\!\ll\!\Phi_{\rm c}$ to hold for CHIME.
Denote by $N_{k}$ the number of sources seen to repeat $k$ times. In the energy limited regime ($\Phi_{\rm lim}\!>\!\Phi_{\rm c}$), $N_{k}\!\propto\lambda_{\rm lim}^{k}r^{3}\!\propto(\Phi_{\rm lim}r^{2})^{k(1-\gamma)}r^{3}\!\propto r^{2k(1-\gamma)+3}$. This shows there is a critical $k_{\gamma}=3/(2(\gamma-1))$ above which the fluence distribution becomes dominated by closer distances (and therefore also lower energies of order $E_{\rm all,k}$), i.e. with a typical distance
@@
$r_{\rm typ}(>\!\Phi;k\!>\!k_{\gamma})\!\approx\!\frac{3}{4}r_{z}\left\{\begin{array}[]{ll}\!1&;\Phi\!<\!\Phi_{\rm all,k}\\ \!\sqrt{\frac{\Phi_{\rm all,k}}{\Phi}}&;\Phi>\Phi_{\rm all,k}.\end{array}\right.$ (6)
@@
Instead, for $k<k_{\gamma}$, if $\Phi<\Phi_{\rm c}$, then energetic bursts with $E_{\rm c}$ can be seen from any $r<r_{z}$, while for $\Phi>\Phi_{\rm c}$ the typical distances are limited by the maximum distance to which $E_{\rm c}$ can be detected
@@
$r_{\rm typ}(>\!\Phi;k\!<\!k_{\gamma})\!\approx\!\frac{3}{4}r_{z}\left\{\begin{array}[]{ll}\!1&;\Phi\!<\!\Phi_{\rm c}\\ \!\sqrt{\frac{\Phi_{\rm c}}{\Phi}}&;\Phi>\Phi_{\rm c}.\end{array}\right.$ (9)
@@
Put simply, assuming a single source rate, very active repeaters will be dominated by nearby sources [^6] In practice, the bottom lines of Eqns. 6, 9 hold only as long as there is at least one source at this distance. We define the typical distance within which there is one source as $n_{\rm src}(4\pi/3)r_{1}^{3}=1\to r_{1}=(3/(4\pi n_{\rm src}))^{1/3}$ (where $n_{\rm src}$ is the source density). In general, sources will reside at $r\!\approx\!\max(r_{\rm typ},r_{1})$., while non-repeaters by distant ones. The fluence distribution of sources repeating $k$ times is $N_{\rm k}(>\Phi)\propto\Phi^{k(1-\gamma)}$. For $k>k_{\gamma}$ the latter is steeper than the result for a constant energy of bursts in Euclidean geometry, $N_{\rm k}(>\Phi)\propto\Phi^{-3/2}$, and therefore it is sub-dominant (the contribution of less energetic $E_{\nu}\approx E_{\rm all,k}$, but closer, sources outpaces that of rarer but more energetic events that are visible to larger distances).
@@
[…]

### v2

@@ Abstract
Do all Fast Radio Burst (FRB) sources repeat? We present evidence that FRB sources follow a Zipf-like distribution, in which the number density of sources is approximately inversely proportional to their burst rate above a fixed energy threshold-even though both the burst rate and number density span many orders of magnitude individually. We introduce a model-independent framework that predicts the distribution of observed fluences and distances, and repetition rates of an FRB population based on an assumed burst rate distribution per source. Using parameters derived directly from observations, this framework simultaneously explains several key features of the FRB population: (i) The observed ratio of repeaters to apparent non-repeaters; (ii) The much lower ratio of apparent non-repeaters to the total number of Soft Gamma Repeater (SGR) sources within the observable Universe; And (iii) the slightly smaller average distances of known repeaters compared to non-repeaters. We further explore how survey parameters, such as radio sensitivity and observation time, influence these statistics. Notably, we find that the fraction of repeaters rises only mildly with improved sensitivity or longer exposure. This weak dependence could be misinterpreted as evidence that not all FRBs repeat. Overall, our results support the idea that a single population-likely magnetars-can account for the full observed diversity of FRB activity, from very inactive FRB sources like SGR 1935+2154 to the most active repeaters.
@@ 1 Introduction
Fast radio bursts (FRBs) exhibit a striking diversity in their observable properties. The isotropic-equivalent energies of extragalactic FRBs with known redshifts span over nine orders of magnitude - from as low as $\sim 6\times 10^{23}\mbox{erg Hz}^{-1}$ in FRB 20200120E (Nimmo et al., 2023) to as high as $\gtrsim 6\times 10^{32}\mbox{erg Hz}^{-1}$ in FRB 20220610A (Ryder et al., 2023). Even individual repeaters can show large variability: for example, bursts from FRB 20201124A cover about five orders of magnitude in energy (Zhang et al., 2022; Kumar et al., 2022). FRB sources also differ greatly in their apparent activity rate (Margalit et al., 2020). Some, like FRBs 20121102A (aka R1) and 20201124A, have been observed to emit dozens of bursts in a single hour (Li et al., 2021; Zhang et al., 2022) while others - such as the Galactic magnetar SGR 1935+2154 - have only produced a single energetic burst with in an observed period of over several years. As discussed in §3, this contrast becomes even more pronounced when comparing sources at similar energies, due to the steep energy dependence of burst rates. A third axis of variation is the inferred number density of sources. For instance, Lu et al. (2022) showed that the volumetric density of highly active extragalactic FRBs like 20180916B and 20121102A is orders of magnitude lower than that of low-activity sources such as SGR 1935+2154 or FRB 20200120E in the nearby M81 galaxy. Given these dramatic differences - in energy output, repetition rate, and source density - it remains an open question whether all FRBs can be unified under a single physical framework, and if so, what underlying parameters account for their apparent diversity.
@@
Perhaps the most striking of the apparent dichotomies, is the observational separation between sources seen only once thus far (“non-repeaters") and those seen multiple times (“repeaters"). The most reliable way to estimate the intrinsic fraction of repeating sources is to consider a blind survey mission such as CHIME. In the first 20 months of its operation, CHIME detected 46 repeating sources (and 14 more candidates), consisting a fraction of $2.6_{-2.59}^{+2.9}\%$ of sources in their sample (Chime/Frb Collaboration et al., 2023). The lack of bimodality in observed burst rates per source between the population of repeaters and (as of yet) non-repeaters detected by CHIME (with observed rates varying continuously from a few per hour to $\lesssim 10^{-3}\mbox{ hr}^{-1}$, Chime/Frb Collaboration et al. 2023), suggests all FRB sources might repeat, and that the low observed fraction is just a reflection of the current sensitivity. The finding that repeaters in this sample have, on average, slightly lower extragalactic DMs than non-repeaters ($\sim 430\mbox{ pc cm}^{-3}$ as compared with $\sim 600\mbox{ pc cm}^{-3}$), could simply be a selection effect - repeaters are by the nature of their identified repetition, likely to be slightly closer than non-repeaters. This does not require intrinsically distinct populations. Similarly, while there are some small differences between repeater and non-repeater host galaxies (the former being slightly less massive, and having lower metallicity, Sharma et al. 2024), these differences are not statistically meaningful and might in addition be subject to further selection biases.
@@
A different way to assess the possibility of a shared origin of repeaters and non-repeaters, is to monitor over a long time known, highly active, repeater sources. The energy distribution per source is found (at least by some studies) to be consistent with the energy distribution of non-repeaters (Wang & Yu, 2017; Zhang et al., 2021). Kirsten et al. (2024) have observed FRB 20201124A for over 2000 hours, using 25-32 m class telescopes. They have found that the slope of the burst energy distribution varies both as a function of time (sporadically) and energy (becoming shallower at higher energies). The higher energy slope is consistent with that of non-repeating sources and the time variation helps explain the apparent large range of repeating rates between sources. Similar results (utilizing 1500 hours of observations) were found for another repeater, FRB 20220912A (Ould-Boukattine et al., 2024).
In this paper, we set out to explore whether the activity rates, source densities, and energies of FRB repeaters / non-repeaters can originate from the same underlying population. Our approach is primarily analytic (backed by numerical calculations) and is independent of the nature of FRB sources. We begin, in §2 by exploring the simplest hypothesis supporting a unified FRB origin: that there is a single source type powering all observed FRBs. Then, in §3 (and §C) we expand the model to allow for variable intrinsic rates and typical burst energies. The results are then compared with observations in §4 and implications for magnetar sources are discussed in §5. We conclude in §6.
@@ 2 Single source rate
for $E_{\rm min}<E_{\nu}<E_{\rm c}$ and $\gamma>1$, where $E_{\nu}$ is the energy per unit frequency, $E_{\rm c}$ is the maximum burst energy per source [^3] Some authors have fit the burst distribution with a functional form of a Schechter function: $d\mathcal{R}/dE_{\nu}\propto(E_{\nu}/E_{\rm c})^{-\gamma}e^{-E_{\nu}/E_{\rm c}}$. In that case $E_{\rm c}$ is the energy marking the transition from the PL distribution to the exponential cut-off. and, in anticipation of §3, we use two subscripts for the normalization rate, $\mathcal{R}_{\rm 0,*}$, the first referring to the activity level of the source (0 denoting the most common and inactive sources or the only sources in the case of a single source activity level explored in this section), and the second to the burst energy at which the normalization is taken. $E_{*}$ is an energy used for the normalization of the rate per source. Throughout the majority of this work we choose $E_{\rm*}=E_{\rm c}$ (an exception to this is discussed in §C).
@@
At the simplest level, a large sky FRB survey can be described by its limiting (spectral) fluence $\Phi_{\rm lim}$ and by the time, $T$, it spends observing each point on the sky  [^4] We first consider a uniform $T$ value across the sky; we provide a generalized treatment in §A. During the data collection period for the 1st CHIME catalog, the median exposure time was $\sim\!20$ hrs (CHIME/FRB Collaboration et al., 2021), which we adopt as a reference value when using a fixed $T$.. While low energy bursts are more common per source (Eq. 1), such bursts cannot be detected to large distances by a fluence-limited survey. At the same time the number of sources increases rapidly with distance. The result is a competition between the detectability of low energy and close-by sources and high energy far-away sources. For $\gamma<2.5$, Euclidean geometry and a uniform volumetric density of sources  [^5] A verification of our results using all relevant cosmological corrections and not assuming a uniform volumetric source density is presented in §A., the competition is won by the far-away sources (see, e.g. Lu & Piro 2019). That is, most detected bursts are from large cosmological distances. This holds up to a distance $r_{z}$ (roughly corresponding to the proper distance up to $z\approx 2$) beyond which the number of sources increases much more slowly, due to a significant departure from Euclidean geometry (as well as a likely intrinsic reduction in the source density, at eras before the peak of star formation history). The combination of $r_{z}$ and $\Phi_{\rm lim}$ yields the typical energy of detectable non-repeater bursts $E_{\rm lim}\equiv 4\pi r_{z}^{2}\Phi_{\rm lim}$ (those that are close to the threshold fluence and arise from distant sources). However, the conclusion that most detected bursts have an energy of order $E_{\rm lim}$ does not necessarily hold for active repeaters, observed to repeat $k\gg 1$ times during the survey’s operation. This is because, if most bursts from sources seen $k$ times have such an energy, then this corresponds to a single mean number of repetitions per source during the observed period, $\lambda_{{\rm lim}}\equiv\lambda_{\rm E_{\rm lim}}=\mathcal{R}(>E_{\rm lim})T$. Assuming Poisson statistics, the probability that a source will repeat $k$ times during $T$ is $P(k)=\lambda^{k}e^{-\lambda}/k!$  [^6] Several studies reported non-Poissonian burst arrival time statistics from prolific repeaters, with bursts tending to cluster in time (Oppermann et al., 2018; Wang et al., 2018; Oostrum et al., 2020). However, Cruces et al. (2021); Jahns et al. (2023) showed that once short separations (on the order of tens of ms, likely corresponding to non-independent sub-bursts) are excluded and periodically modulated activity is accounted for, the remaining distribution is consistent with a Poisson process. Since deviations from Poissonian behavior would require additional model parameters, we adopt a Poissonian model here to avoid overfitting. This assumption can be revisited if future observations provide strong evidence for non-Poissonian statistics.. Therefore, $\lambda/2$ is approximately the ratio between repeaters (sources observed at least twice) and non-repeaters and calibration with the observed ratio leads to $\lambda_{\rm lim}\approx 0.05$. $P(k)$ then becomes vanishingly small for large $k$. For example $P(k\geq 10)/P(k\geq 1)\approx\lambda^{10}/10!\sim 4\times 10^{-20}$. The fact that CHIME has detected multiple sources with more than 10 repetitions clearly means that such a picture is too simplistic and that the distance and / or intrinsic rate per source must be different for sources detected multiple times. The first point is addressed below and the second in §3.
Under the assumption of a single source type, the value of $\lambda_{\rm lim}\approx 0.05$ calibrated by the repeater / non-repeater ratio can be used to deduce the rate per source above $E_{\rm lim}$,
@@
$\mathcal{R}(>E_{\rm lim})\!\approx\!2.5\times 10^{-3}\left(\frac{20\mbox{ hr}}{T}\right)\left(\frac{\lambda_{\rm lim}}{0.05}\right)\mbox{ hr}^{-1}$
@@
[…]
@@
where $\lambda_{\rm c}\!\equiv\!\lambda_{E_{\rm c}}$ and (since by definition $k\!\geq\!1$) the last inequality necessarily holds for $\lambda_{\rm lim}\!<\!1$.
@@
We define the fluences $\Phi_{\rm c}\!=\!E_{\rm c}/(4\pi r_{z}^{2})$ and $\Phi_{\rm all,k}\!=\!E_{\rm all,k}/(4\pi r_{z}^{2})\!\approx\!(\lambda_{\rm c}/k)^{1\over\gamma-1}\Phi_{\rm c}$ which are respectively the fluences corresponding to bursts with energy $E_{\rm c}$ and $E_{\rm all,k}$ as viewed from a distance $r_{z}$. Below $\Phi_{\rm all,k}$, roughly all sources in the Universe are observed by the survey $\geq k$ times. Taking $E_{\rm c}\!\approx\!3\!\times\!10^{33}\mbox{erg Hz}^{-1},\gamma\!\approx\!1.7$ as inferred from FRB energy distribution studies calibrated by observations (Lu & Piro, 2019; James et al., 2022; Hashimoto et al., 2022; Shin et al., 2023; Lin & Zou, 2024; Gupta et al., 2025) and recalling that $\lambda_{\rm c}=\lambda_{\rm lim}(E_{\rm c}/E_{\rm lim})^{1-\gamma}$ (Plugging $E_{\nu}=E_{\rm c}$ in Eq. 1 and multiplying by $T$), we find $\lambda_{\rm c}\!\approx\!7\!\times\!10^{-3},\Phi_{\rm c}\!\approx\!100\mbox{Jy ms},\Phi_{\rm all,1}\!\approx\!0.07\mbox{Jy ms}$, so we expect the ordering $\Phi_{\rm all,1}\!\ll\!\Phi_{\rm lim}\!\ll\!\Phi_{\rm c}$ to hold for CHIME.
Denote by $N_{k}$ the number of sources seen to repeat $k$ times above a fluence $\Phi_{\rm lim}$. In the energy limited regime ($\Phi_{\rm lim}\!>\!\Phi_{\rm c}$), $N_{k}\!\propto\lambda_{\rm lim}^{k}r^{3}\!\propto(\Phi_{\rm lim}r^{2})^{k(1-\gamma)}r^{3}\!\propto r^{2k(1-\gamma)+3}\propto E^{k(1-\gamma)+3/2}$ (where in the last transition we have used $E\propto\Phi_{\rm lim}r^{2}$). This shows there is a critical $k_{\gamma}=3/(2(\gamma-1))$ and that the observed distribution of distances / energies will depend on $k$ relative to $k_{\gamma}$. Consider first $k\!>\!k_{\gamma}$. In this range, $N_{k}$ decreases with $E$, and bursts with lower energy are detected more commonly than those with larger energies. This holds down to the energy $E_{\rm all,k}$, below which all sources are detected at least $k$ times. Therefore at $\Phi<\Phi_{\rm all,k}$, all sources are observable. For a uniform density of sources, and Euclidean geometry, their cumulative distance distribution is $N_{k}(r)=(r/r_{z})^{3}$ and the median distance is $r=0.5^{1/3}r_{z}\approx 0.8r_{z}$. At larger fluences, the most distant sources are missed. Fixing the energy of the bursts at $E_{\rm all,k}$ as explained above, we see that the maximum detectable distance becomes $r(\Phi)=(E_{\rm all,k}/(4\pi\Phi))^{1/2}=r_{z}(\Phi_{\rm all,k}/\Phi)^{1/2}$. Overall,
@@
$r_{\rm typ}(>\!\Phi;k\!>\!k_{\gamma})\!\approx\!0.8r_{z}\left\{\begin{array}[]{ll}\!1&;\Phi\!<\!\Phi_{\rm all,k}\\ \!\sqrt{\frac{\Phi_{\rm all,k}}{\Phi}}&;\Phi>\Phi_{\rm all,k}.\end{array}\right$. (6)
@@
Instead, for $k<k_{\gamma}$, the most energetic bursts, with $E_{\nu}\approx E_{\rm c}$ are the ones detected most often. If $\Phi<\Phi_{\rm c}$, then energetic bursts with $E_{\rm c}$ can be seen from any $r<r_{z}$, while for $\Phi>\Phi_{\rm c}$ the typical distances are limited by the maximum distance to which $E_{\rm c}$ can be detected
@@
$r_{\rm typ}(>\!\Phi;k\!<\!k_{\gamma})\!\approx\!0.8r_{z}\left\{\begin{array}[]{ll}\!1&;\Phi\!<\!\Phi_{\rm c}\\ \!\sqrt{\frac{\Phi_{\rm c}}{\Phi}}&;\Phi>\Phi_{\rm c}.\end{array}\right$. (9)
@@
Put simply, assuming a single source rate and for $\Phi_{\rm all,k}<\Phi<\Phi_{\rm c}$, very active repeaters will be dominated by nearby sources (bottom part of Eq. 6) [^7] In practice, the bottom lines of Eqns. 6, 9 hold only as long as there is at least one source at this distance. We define the typical distance within which there is one source as $n_{\rm src}(4\pi/3)r_{1}^{3}=1\to r_{1}=(3/(4\pi n_{\rm src}))^{1/3}$ (where $n_{\rm src}$ is the source density). In general, sources will reside at $r\!\approx\!\max(r_{\rm typ},r_{1})$., while non-repeaters by distant ones (top part of Eq. 9). The fluence distribution of sources repeating $k$ times is $N_{\rm k}(>\Phi)\propto\Phi^{k(1-\gamma)}$. For $k>k_{\gamma}$ the latter is steeper than the result for a constant energy of bursts in Euclidean geometry, $N_{\rm k}(>\Phi)\propto\Phi^{-3/2}$, and therefore it is sub-dominant (the contribution of less energetic $E_{\nu}\approx E_{\rm all,k}$, but closer, sources outpaces that of rarer but more energetic events that are visible to larger distances).
@@
$N_{\geq k}(>\!\Phi)\!=\!\sum_{k}^{\infty}\int_{0}^{\min[r_{z},\sqrt{\frac{E_{\rm c}}{4\pi\Phi}}]}dr\frac{e^{-\lambda_{E}}\lambda_{E}^{k}}{k!}n_{\rm src}4\pi r^{2}$